# HEPHY-ML-25-02 BitHEP — The Limits of Low-Precision ML in HEP Claudius Krause1, Daohan Wang1 and Ramon Winterhalder2

### v1

@@ 5.1 Normalizing Flow — CaloINN
5.2 Conditional Flow Matching — CaloDREAM
@@
In detail, CaloDREAM consists of two networks, both trained with conditional flow matching: (i) an autoregressive transformer, called the energy network, to learn the 45 layer energies, and (ii) a vision transformer, called the shape network, learning the normalized showers. The self-attention mechanism in the transformer layers is highly beneficial for modeling the sparsity of calorimeter showers and the correlations across detector locations. Crucially, the introduction of patching—which groups nearby voxels into coarser units—reduces the impact of the quadratic scaling and enables the model to scale efficiently to larger calorimeter geometries such as DS3. At the same time, conditional flow matching ensures efficient training of the underlying continuous normalizing flow model.
@@
Quantizes the core elements of the ViT blocks—QKV matrices, projections, and MLPs—while keeping the embedding networks and the final MLP layer unquantized. This configuration results in 63.8% of the shape network being quantized.
@@ 6 Conclusions and Outlook
Larger networks can be quantized more easily. Our experiments with different generative architectures used in detector simulation—specifically the normalizing flow-based CaloINN and the flow matching-based CaloDREAM — indicate that larger models often exhibit smoother quantization behavior. For instance, with around 66% of its parameters quantized, CaloINN experiences a noticeable drop in sample quality, whereas CaloDREAM, which also quantizes roughly 63.8% of its shape network, exhibits only minor performance degradation. These observations underscore how larger networks provide greater representational capacity, making them more resilient to the information loss introduced by low-bit representations. Consequently, these results highlight the potential of quantized, large-scale generative models in tackling the complex, high-dimensional tasks typical of modern particle physics experiments.
Performance depends on the layer chosen for quantization. In our generative studies for detector simulation, CaloINN benefits from selectively quantizing all layers within the central bijectors (BlockCentral), thereby maintaining decent sample quality even at about 66% quantized weights, while fully quantizing all layers (99.9%) severely degrades performance. Likewise, in CaloDREAM, quantizing the elements within the shape network’s ViT blocks (QKV matrices, projections, and MLPs) results in minimal performance loss; however, once the embedding layers are quantized, performance drops considerably. These findings underscore how carefully selecting which sections of the network to quantize—central layers versus outer layers, embedding networks versus projection layers—can preserve model fidelity and better adapt to each architecture’s demands, especially for high-dimensional tasks in particle physics. Further studies in automatic, heterogeneous quantization [201, 202] are needed to fully extract the best performance of the networks.
Quantizing more layers generally degrades model performance. Our SMEFTNet experiments demonstrate that as the proportion of quantized layers increases, the precision of the predictions decreases. The fully quantized SMEFTNet-Bit shows the largest degradation, while selectively quantizing only 30% of the linear layers (within the message-passing block) leads to a minimal drop in performance. These results highlight the trade-off between model compression and accuracy, emphasizing the need for selective quantization to balance efficiency and precision.

### v2

@@ 5.1 Normalizing Flow – CaloINN
| Dataset | Operations | Default | NNCentral | All | ||
| ds1– | FLOPs | |||||
| SignOPs | - | |||||
| IntOPs | - | |||||
| Total FLOPs | (92%) | (6%) | ||||
| ds1–+ | FLOPs | |||||
| SignOPs | - | |||||
| IntOPs | - | |||||
| Total FLOPs | (94%) | (6%) | ||||
| ds2 | FLOPs | |||||
| SignOPs | - | |||||
| IntOPs | - | |||||
| Total FLOPs | (99.7%) | (5.6%) | ||||
@@
To estimate the resource consumption of the quantized CaloINN, we compute the number of FLOPs, IntOPs, and SignOPs of the three different types of coupling layers that make up our quantization strategies in Tab 6. While these number take into account the batchnormalizations in the NN, they do not include the FLOPs that are used in the spline evaluation.
5.2 Conditional Flow Matching – CaloDREAM
@@
In detail, CaloDREAM consists of two networks, both trained with conditional flow matching: (i) an autoregressive transformer, called the energy network, to learn the 45 layer energies, and (ii) a vision transformer, called the shape network, learning the normalized showers. The self-attention mechanism in the transformer layers is highly beneficial for modeling the sparsity of calorimeter showers and the correlations across detector locations. Crucially, the introduction of patching, which groups nearby voxels into coarser units, reduces the impact of the quadratic scaling and enables the model to scale efficiently to larger calorimeter geometries such as DS3. At the same time, conditional flow matching ensures efficient training of the underlying continuous normalizing flow model.
@@
Quantizes the core elements of the ViT blocks, i.e. QKV matrices, projections, and MLPs, while keeping the embedding networks and the final MLP layer unquantized. This configuration results in 63.8% of the shape network being quantized.
@@
| Block | Operations | Regular | Quantized | |
| Energy Network | FLOPs | |||
| SignOPs | - | |||
| IntOPs | - | |||
| Total FLOPs | (10%) | |||
@@
In tables 8 and 9 we give an estimate of the computational resources required for different components of CaloDREAM. The numbers are based on a single run through the networks, so they only show a fraction of the FLOPs required to generate a full shower. In addition, we also do not include the transformer decoder and encoder networks of the energy network, as they are not quantized in our studies.
| Block | Operations | Regular | No embedding | Full | ||
| Embedding Layers | FLOPs | |||||
| SignOPs | - | - | ||||
| IntOPs | - | - | ||||
| Total FLOPs | (100%) | (6%) | ||||
| Diffusion Transformer Blocks | FLOPs | |||||
| SignOPs | - | |||||
| IntOPs | - | |||||
| Total FLOPs | (38%) | (38%) | ||||
@@ 6 Conclusions and Outlook
Larger networks can be quantized more easily. Our experiments with different generative architectures used in detector simulation, specifically the normalizing flow-based CaloINN and the flow matching-based CaloDREAM, indicate that larger models often exhibit smoother quantization behavior. For instance, with around 66% of its parameters quantized, CaloINN experiences a noticeable drop in sample quality, whereas CaloDREAM, which also quantizes roughly 63.8% of its shape network, exhibits only minor performance degradation. These observations underscore how larger networks provide greater representational capacity, making them more resilient to the information loss introduced by low-bit representations. Consequently, these results highlight the potential of quantized, large-scale generative models in tackling the complex, high-dimensional tasks typical of modern particle physics experiments.
Performance depends on the layer chosen for quantization. In our generative studies for detector simulation, CaloINN benefits from selectively quantizing all layers within the central bijectors (BlockCentral), thereby maintaining decent sample quality even at about 66% quantized weights, while fully quantizing all layers (99.9%) severely degrades performance. Likewise, in CaloDREAM, quantizing the elements within the shape network’s ViT blocks (QKV matrices, projections, and MLPs) results in minimal performance loss; however, once the embedding layers are quantized, performance drops considerably. These findings underscore how carefully selecting which sections of the network to quantize – central layers versus outer layers, embedding networks versus projection layers – can preserve model fidelity and better adapt to each architecture’s demands, especially for high-dimensional tasks in particle physics. Further studies in automatic, heterogeneous quantization [202, 203] are needed to fully extract the best performance of the networks.
Fully quantized models show the largest performance degradation. Our SMEFTNet experiments demonstrate that the fully quantized variant performs worst among all tested configurations. Specifically, SMEFTNet-Bit100 exhibits a larger degradation in accuracy than the configurations where only a subset of layers is quantized (SMEFTNet-Bit30 or SMEFTNet-Bit70). These results highlight the trade-off between model compression and accuracy, emphasizing the need for selective quantization to balance efficiency and precision.